40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3824     Entry time: Fri Oct 29 14:16:26 2010
Author: Jenne 
Type: Update 
Category: SUS 
Subject: PRM baking 

[Suresh, Jenne]

We took a look-see at the PRM after the gluing from last night.  The balance is still okay.  The reflected beam is a teeny bit below the laser aperture (center of the beam maybe ~2mm below, so ~1mRad low).  This is within our okay range, since the DC offset that the OSEMs will give will be even more, and the coils can definitely handle this kind of offset.

We took the optic out of the tower, and gave it to Bob and Daphen to bake over the weekend.

ELOG V3.1.3-